Grade 11General Physics

two forces are in the ratio 1:2 named p and q.if their resultant is at an angle tan inverse root 3 /2 to vector p then the angle between p and q is ?

Prachi
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
p = F (let)
q = 2F (let)
R = root(F^2 + 4F^2 + 4F^2costhetha)
 
cos @ = F/R = > F/F root(5 + 4costhetha) = > 1/root(5 + 4costhetha)
cos(arctanroot(3/2)) =  1/root(5 + 4costhetha)
or
cos(arccos(root(2/5) =  1/root(5 + 4costhetha)
root(2/5) =  1/root(5 + 4costhetha)
squaring both sides,
10 + 8costhetha = 5
8costhetha  =  -5
costhetha = -5/8
thetha = arccos(-5/8)